# Rigidity of the escaping set of polynomial automorphisms of $\mathbb{C}^{2}$

### v1

@@ 1. Introduction
is of the form $f=L\circ H^{s}\text{ with }s\in\mathbb{Z},$ where $L$ is an affine map that preserves both $K^{\pm}$. In addition,
@@
Here, the assumption that $f(K^{-})=K^{-}$, which was imposed in [5] is not required, and this is the main thrust of this theorem. The proof consists of several steps that are summarized as follows.
@@ 2. An intermediate cover of the escaping set $U^{+}$
This is a generalization of the results in [15, Section 8]. We will briefly revisit the steps, mostly following [17, Chapter 7]. Note that
@@ 4.2. The group ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$
Thus, $f$ preserves both $K^{\pm}$. Hence by Theorem 1.1 in [5], $f=L\circ H^{s^{\prime}}$ where $L$ is an affine map that preserves both $K^{\pm}$. ∎
@@
Let $H$ be a generalised Hénon map of the form (1.1) and $\phi\in{\rm Aut}(\mathbb{C}^{2})$ such that $\phi(K_{H}^{+})\subset K_{H}^{+}$. Then $\phi=L\circ H^{s}$ where $L$ is a linear map preserving both $K_{H}^{\pm}$ and $s$ is an integer.
@@
Finally, we use the above to prove a version of [9, Theorem A] under comparatively weaker assumptions.
@@
Note that, $\phi^{-1}\circ H\circ\phi(K_{F}^{+})\subset K_{F}^{+}$ and is also an automorphism of $\mathbb{C}^{2}$. So by the above result, there exist an affine map $L$ and $s\in\mathbb{Z}$ such that $\phi^{-1}\circ H\circ\phi=L\circ F^{s}$. In particular,
$L^{-1}\circ\phi^{-1}\circ H\circ\phi\circ L=F^{s}\circ L,\text{ i.e., }(\phi\circ L)^{-1}\circ H\circ(\phi\circ L)=F^{s}\circ L.$
@@
$(\phi\circ L)^{-1}\circ H^{n_{k}}\circ(\phi\circ L)=F^{sn_{k}}\circ L^{n_{0}}.$
@@
Thus $f\in{\rm Aut}(\mathbb{C}^{2})$ as $\mathbb{C}^{2}=\cup_{n}\Omega_{cd^{n}}$, with $f(K^{+})=K^{+}$. Thus by Theorem 1.1, $f=L\circ H^{s}$ where $s\in\mathbb{Z}$ where $L$ is an affine map preserving $K^{\pm}$. Further $f(\Omega_{c})=\Omega_{c}$, hence $f=L$. ∎
Finally, we conclude with the following application on biholomorphic Short $\mathbb{C}^{2}$’s
@@
Let $H$ and $F$ be generalised Hénon maps of the form (1.1) of the same degree $d\geq 2$ and assume that $d=\mathbf{p}^{m}$ for some prime $\mathbf{p}$ and $m\geq 1$. Suppose that $\Omega_{H,c}$ and $\Omega_{F,c^{\prime}}$ are biholomorphic, where
@@
for some $c,c^{\prime}>0$. Then there exist affine maps $A_{1}$ and $A_{2}$ such that
$H=A_{1}\circ F\circ A_{2}.$
@@
Thus $\phi$ extends to an automorphism of $\mathbb{C}^{2}$ such that $\phi(K_{F}^{+})=K_{H}^{+}$, or equivalently their corresponding escaping sets $U_{F}^{+}$ and $U_{H}^{+}$ are biholomorphic, via the map $\phi$. In particular by Theorem 1.1, each $\mathfrak{a}_{n}$, $n\geq 1$ is an affine map such that $\mathfrak{a}_{n}(K_{H}^{-})=K_{H}^{-}$, $G_{H}^{-}\circ{\mathfrak{a}}_{n}=G_{H}^{-}$ and $\mathfrak{a}_{n}\in\mathcal{L}$. Since there are only finitely many choice for every $\mathfrak{a}_{n}$, it follows from (5.2) that $\phi(K_{H}^{-})=K_{F}^{-}$ or $\phi(U_{H}^{-})=U_{F}^{-}$.
Thus $\phi^{*}(\mu_{H}^{-})$ and $\phi^{*}(\mu_{H}^{+})$ are $dd^{c}$-closed $(1,1)$-positive currents supported on $K_{H}^{-}$ and $K_{H}^{-}$, respectively. By arguing similarly as in the proof of Theorem A in [9, page 3748], there exist constants $a,b>0$ such that
@@
Hence $b=a^{-1}$.
@@
Hence $\|f(p)\|<M\|p\|+C$ for every $p\in\mathbb{C}$. This proves that $f$ is an affine map. Further, from (5.2), if $s\leq 0$,
$\mathfrak{a}_{-s+1}=\phi^{-1}\circ F^{-s+1}\circ\phi\circ H^{s-1}\text{ and }F^{s}\circ\phi=\phi\circ H^{s}\circ\mathfrak{a}_{-s}^{-1}.$
Thus $F^{s}\circ\phi$ is an affine map. Finally,
$H=\left(F^{s}\circ\phi\right)^{-1}\circ F\circ\left(\phi\circ H^{s}\circ\mathfrak{a}_{-s+1}^{-1}\right),$
which completes the proof. Since $\phi$ extends to an automorphism, (5.2) is true for $n\leq 0$. Hence, a similar argument works if $s>0$. ∎
@@
Note that the isomorphism induced by the map $\phi$ (in the above proof) on the fundamental group $\mathbb{Z}\left[\frac{1}{d}\right]$, is $\mathfrak{i}_{\phi}(x)=u_{\phi}x$, where $u_{\phi}$ is an unit in the ring $\mathbb{Z}\left[\frac{1}{d}\right]$. Suppose $u_{\phi}d^{n}\neq 1$ for every $n\in\mathbb{Z}$, then the above argument will not work, hence Theorem 5.1 could not be stated generally for arbitrary degrees $d\geq 2$.

### v2

# arXiv:2601.07681v2[math.CV] 01 Mar 2026

@@ 1. Introduction
is of the form $f^{r}=L\circ H^{s}\text{ with }r,s\in\mathbb{Z}$ and $r>0$, where $L$ is an affine map that preserves both $K^{\pm}$. In addition,
@@
Here, $f$ is essentially a rational root of the Hénon map $H$, in the sense of [9]. However, the assumption that $f(K^{-})=K^{-}$, which was imposed in [5] is not required, and this is the main thrust of this theorem. Further, in the spirit of [5, Theorem 1.1] it follows that there exists a Hénon map $\rho_{H}$ (an irreducible root of $H$, again in the sense of [9]) such that
$f=L\circ\rho_{H}^{s}\text{ and }f\circ H=\tilde{L}\circ H\circ f$
where $L,\tilde{L}\in\mathcal{L}$ and $s\in\mathbb{Z}$. The proof consists of several steps that are summarized as follows.
@@ 2. An intermediate cover of the escaping set $U^{+}$
This is a generalization of the results in [16, Section 8]. We will briefly revisit the steps, mostly following [18, Chapter 7]. Note that
@@ 4.2. The group ${\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$
Thus, $f$ preserves both $K^{\pm}$. Hence by Theorem 1.1 in [5], $f=L\circ\rho_{H}^{s^{\prime}}$ where $L$ is an affine map that preserves both $K^{\pm}$ and $\rho_{H}$ is a Hénon map preserving $K_{H}^{\pm}$ such that degree of $\rho_{H}$ (say $r^{\prime}$) divides the degree of $f$. Also by Theorem 1.1 in [5], $r^{\prime}$ divides $d$. Thus there exist positive integer $r$ such that degree of $f^{r}$ is divisible by $d$. Further, there exits an integer $s$ such that $L=f^{r}\circ H^{-s}\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$, i.e., by Theorem 4.7, $L$ is an affine map. This completes the proof. ∎
@@
Let $H$ be a generalised Hénon map of the form (1.1) and $\phi\in{\rm Aut}(\mathbb{C}^{2})$ such that $\phi(K_{H}^{+})\subset K_{H}^{+}$. Then $\phi^{r}=L\circ H^{s}$ where $L$ is a linear map preserving both $K_{H}^{\pm}$ and $r,s$ are integers.
@@
Finally, we use the above to prove a version of [10, Theorem A] under comparatively weaker assumptions.
@@
Note that, $\phi^{-1}\circ H\circ\phi(K_{F}^{+})\subset K_{F}^{+}$ and is also an automorphism of $\mathbb{C}^{2}$. So by the above result, there exist an affine map $L$ and $r,s\in\mathbb{Z}$ with $r>0$ such that $\phi^{-1}\circ H^{r}\circ\phi=L\circ F^{s}$. In particular,
$L^{-1}\circ\phi^{-1}\circ H^{r}\circ\phi\circ L=F^{s}\circ L,\text{ i.e., }(\phi\circ L)^{-1}\circ H^{r}\circ(\phi\circ L)=F^{s}\circ L.$
@@
$(\phi\circ L)^{-1}\circ H^{rn_{k}}\circ(\phi\circ L)=F^{sn_{k}}\circ L^{n_{0}}.$
@@
Thus $f\in{\rm Aut}(\mathbb{C}^{2})$ as $\mathbb{C}^{2}=\cup_{n}\Omega_{cd^{n}}$, with $f(K^{+})=K^{+}$. Thus by Theorem 1.1, $f^{r}=L\circ H^{s}$ where $r,s\in\mathbb{Z}$ where $L$ is an affine map preserving $K^{\pm}$. Further $f^{r}(\Omega_{c})=\Omega_{c}$, hence $f^{r}=L$ or $f\in{\rm Aut}_{1}(U^{+})\cap{\rm Aut}(\mathbb{C}^{2})$. Now the proof follows from Theorem 4.7. ∎
Finally, we conclude with the following applications on biholomorphic Short $\mathbb{C}^{2}$’s
@@
Let $H$ and $F$ be generalised Hénon maps of the form (1.1) of the same degree $d\geq 2$. Suppose that $\Omega_{H,c}$ and $\Omega_{F,c^{\prime}}$ are biholomorphic, where
@@
for some $c,c^{\prime}>0$. Then there exist polynomial maps $P_{1}$ and $P_{2}$ such that
$H=P_{1}\circ F\circ P_{2}.$
In particular, if $d=\mathbf{p}$ for some prime $\mathbf{p}\geq 2$ then $P_{1}$ and $P_{2}$ are affine maps.
@@
Thus $\phi$ extends to an automorphism of $\mathbb{C}^{2}$ such that $\phi(K_{F}^{+})=K_{H}^{+}$, or equivalently their corresponding escaping sets $U_{F}^{+}$ and $U_{H}^{+}$ are biholomorphic, via the map $\phi$. In particular by Theorem 1.1, each $\mathfrak{a}_{n}$, $n\geq 1$ is an affine map such that $\mathfrak{a}_{n}(K_{H}^{-})=K_{H}^{-}$, $G_{H}^{-}\circ{\mathfrak{a}}_{n}=G_{H}^{-}$ and $\mathfrak{a}_{n}\in\mathcal{L}$. Since there are only finitely many choice for every $\mathfrak{a}_{n}$, it follows from (5.2) that $\phi(K_{H}^{-})=K_{F}^{-}$ or $\phi(U_{H}^{-})=U_{F}^{-}$. Thus the statement follows from Corollary 4.9.
Now $\phi^{*}(\mu_{H}^{-})$ and $\phi^{*}(\mu_{H}^{+})$ are $dd^{c}$-closed $(1,1)$-positive currents supported on $K_{H}^{-}$ and $K_{H}^{-}$, respectively. By arguing similarly as in the proof of Theorem A in [10, page 3748], there exist constants $a,b>0$ such that
@@
Hence $b=a^{-1}$. Assuming $d=\mathbf{p}$ we prove the following
@@
Hence $\|f(p)\|<M\|p\|+C$ for every $p\in\mathbb{C}$. This proves that $f$ is an affine map. Similarly $g=F^{-1}\circ f\circ H$ is also an affine map as it induces identity on the fundamental group. Thus
$F=f\circ H\circ g.\qed$
Corollary 5.2.
Let $F$ be generalised Hénon maps of the form (1.1) and $H$ be a simple Hénon map of degree $d_{H}=\mathbf{p}$ for some prime $\mathbf{p}\geq 2$. Suppose that $\Omega_{H,c}$ and $\Omega_{F,c^{\prime}}$ are biholomorphic then there exist $m\geq 1$ and affine maps $A_{1}$ and $A_{2}$ such that
$H^{m}=A_{1}\circ F\circ A_{2}.$
Proof.
As $\Omega_{H,c}$ and $\Omega_{F,c^{\prime}}$ are biholomorphic,
$\mathbb{Z}\left[\frac{1}{\mathbf{p}}\right]\simeq\uppi_{1}\left(\Omega_{H,c}^{\prime}\right)\simeq\uppi_{1}\left(\Omega_{F,{c^{\prime}}}^{\prime}\right).$
Hence $d_{F}=\mathbf{p}^{m}$ for some $m\geq 1$ and the degrees of $F$ and $H^{m}$ are same. Let $H^{\prime}=H^{m}$, then $G_{H}=G_{H^{\prime}}$. Now by replicating the above proof with $H=H^{\prime}$ (except, the map $f=\phi\circ H^{s}$) proves the statement. ∎
@@
The above Corollary 5.2 proves that there exists a positive integer $m\geq 1$ such that $m$-the root of an appropriate linear conjugate of $F$, which is known to be a polynomial by [9] is $H$, provided they induce biholomorphic short $\mathbb{C}^{2}$’s at certain values.
Finally also note, the isomorphism induced by the map $\phi$ (in the above proofs) on the fundamental group $\mathbb{Z}\left[\frac{1}{d}\right]$, is $\mathfrak{i}_{\phi}(x)=u_{\phi}x$, where $u_{\phi}$ is an unit in the ring $\mathbb{Z}\left[\frac{1}{d}\right]$. Suppose $u_{\phi}d^{n}\neq 1$ for every $n\in\mathbb{Z}$, then the above argument will not work, hence the above results (on affine conjugation) could not be stated generally for arbitrary degrees $d\geq 2$.